# AI-Assisted Game Management Decisions: A Fuzzy Logic Approach to Real-Time Soccer Substitutions

### v3

@@ Abstract
In elite soccer, substitution decisions entail significant financial and sporting consequences yet remain heavily reliant on intuition or predictive models that merely mimic historical biases. This paper introduces a Fuzzy Logic-based Decision Support System (DSS) designed for real-time, prescriptive game management. Unlike traditional Machine Learning approaches that encounter a “predictive ceiling” by attempting to replicate human behavior, our system audits performance through an objective, rule-based inference engine. We propose a methodological advancement by reformulating the PlayeRank metric into a Cumulative Mean with Role-Aware Normalization, eliminating the play-time exposure bias inherent in cumulative sum models to enable accurate intra-match comparison. The system integrates this refined metric with physiological proxies (fatigue) and contextual variables (disciplinary risk modulated by tactical role) to calculate a dynamic Substitution Priority ($P_{final}$). Validation via a case study of the 2018 FIFA World Cup match between Brazil and Belgium demonstrates the system’s ecological validity: it not only aligned with expert consensus on executed substitutions (e.g., Gabriel Jesus) but, crucially, identified high-risk scenarios ignored by human decision-makers. Specifically, the model flagged the “Fagner Paradox”, a maximum-priority defensive risk, minutes before a critical yellow card, and detected the “Lukaku Paradox,” where an isolated assist masked a severe drop in participation. These results confirm that Fuzzy Logic offers a transparent, explainable, and superior alternative to “black-box” models for optimizing real-time tactical decisions.
@@ 1 Introduction
The financial stakes in modern soccer have never been higher, creating a vast economic disparity between clubs. In the 2023/24 season, Manchester City earned £175.9 million in prize money for winning the Premier League title GE_PremierLeague. In contrast, Fluminense, a Brazilian club with more modest economic power, received a total of US\$50.71 million for reaching the semifinals of the 2025 Club World Cup—a sum that could have escalated to US\$40 million for winning the finalESPN_Mundial.
Despite these immense financial consequences, one of the most impactful in-game decisions: player substitutions, often remains guided by intuition and limited visible cues. As top coaches like Guardiola acknowledge, the timing and choice of a substitution can be the difference between winning titles and missing crucial opportunities.
This gap is particularly evident because tactical decision-making during a soccer match is one of the most complex and high-impact processes in team management. Substitutions, changes in tactical posture, or defensive adjustments are frequently made by coaches based on a combination of empirical knowledge and a multifactorial assessment of inherently imprecise variables. Factors like ”the player looks tired,” ”performance has recently dropped,” or ”the risk of expulsion is high” are not binary, but rather gradual. Performance evaluation is, in itself, a notable challenge. Although the availability of large-scale event data has grown, there is no single, universally accepted metric that captures all facets of a player’s performance Pappalardo2019.
This work addresses the problem of tactical decision-making under uncertainty through the implementation of a fuzzy logic control system. Fuzzy logic is an artificial intelligence tool particularly suited for this domain, as it allows for modeling ”expert knowledge” (a coach’s rules) and handling vague linguistic concepts, such as ”high age,” ”low performance,” or ”medium fatigue” (Marliere2017). The objective of this paper is to design, implement, and validate a decision support system (DSS) focused specifically on player substitution priority. The system uses four inputs (performance, fatigue, age, and disciplinary risk) to generate a fuzzy output (Substitution Priority), providing a quantitative tool to assist the coaching staff with one of their most critical real-time decisions.
2 Literature Review
@@
The academic literature on soccer analytics provides the theoretical foundation for this work, situated at the intersection of performance evaluation and tactical decision modeling. A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo2019, defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
@@
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by Mohandas et al. mohandas, for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
Despite these advances, a gap remains in the literature. Existing works either focus on performance evaluation (Pappalardo2019; LACCEI2023) or on tactical reasoning frameworks Marliere2017, but there is no decision support system that integrates a validated performance metric with contextual match factors—such as fatigue, momentum shifts, and disciplinary risk—for the specific purpose of informing substitution decisions. This work addresses this gap by combining temporally resolved performance evaluation with fuzzy decision modeling within a unified substitution strategy framework.
3 Methodology

### v5

# Auditing Human Decision-Making in High-Stakes Environments via Prescriptive AI: A Stress-Test on Real-Time Tactical Management

@@ Abstract
High-stakes decision-making is often compromised by cognitive biases and outcome dependency. Current AI models typically mimic historical human behavior, inheriting these biases and limiting their utility for normative improvement. Here, we introduce a Prescriptive AI framework designed to audit, rather than automate, human judgment in real-time environments. By decoupling decision quality from stochastic outcomes, we quantify ”decision latency” and status quo bias in elite soccer management—a high-pressure adversarial domain. Analyzing 2018 FIFA World Cup data, our system exposes critical risk states, such as performance collapse following salient positive events (e.g., an assist), which human experts systematically overlook due to outcome bias. These findings demonstrate that interpretable auditing systems can reveal structural flaws in human reasoning that predictive models obscure. This approach establishes a paradigm for Human-AI interaction prioritizing epistemic accountability over predictive mimicry in safety-critical domains.
@@ 1 Introduction
Decision making under uncertainty in high-stakes environments constitutes a central challenge for human judgment. Settings such as clinical triage, operational risk management, and strategic resource allocation are characterized by irreversible actions, asymmetric consequences, and a critical dependence on timely and well-calibrated decisions. In such domains, decision errors are costly not only because of incorrect outcomes, but because they are often unjustified, delayed, or poorly aligned with underlying risk.
Despite these constraints, computational decision-support systems have been dominated by a fundamentally predictive paradigm. Models are typically optimized to replicate historical decision patterns, achieving statistical accuracy by mimicking prior human behavior. However, reliance on behavioral imitation is problematic: historical decisions often encode cognitive biases and institutional inertia. As a result, predictive systems create a structural mismatch between accuracy and decision quality, reaching a documented “predictive ceiling”—for instance, state-of-the-art substitution models plateau at approximately $70\%$ accuracy by merely cloning human choices (mohandas). Moreover, standard evaluation practices in stochastic environments suffer from outcome bias, judging decision quality by realized results rather than by the epistemic justification available at the time of choice (BaronHershey1988).
Importantly, this work does not aim to improve predictive accuracy, automate decision-making, or replace human expertise. Instead, it focuses on auditing the quality of human judgment under uncertainty by separating epistemic justification from stochastic outcomes. From this perspective, the limitations of predictive systems are not merely due to data insufficiency, but reflect a normative flaw inherent to supervised learning in agentic contexts. By mapping contextual inputs to historical decisions, such systems inevitably reproduce embedded cognitive biases, including status quo bias (reluctance to change strategy) and sunk cost fallacy (persistence in failing courses of action). Consequently, predictive models tend to validate conservative human behavior rather than reveal when deviation is epistemically warranted.
To address this limitation, we distinguish between Predictive AI, which forecasts events or replicates past decisions, and Prescriptive AI, defined here as a class of systems whose primary function is to audit, justify, and support human decision-making under uncertainty. This perspective aligns with decision-theoretic and epistemic accounts of agency, in which rational choice is defined as a coherent state transition at time $t$, independent of the stochastic outcome realized at time $t+n$ (vanBenthem2011). While prescriptive analytics and interpretable decision-support frameworks have been advocated in theory (Bertsimas2020; Lepenioti2020; Rudin2019), empirical demonstrations in realistic, adversarial environments remain scarce. Black-box models are particularly ill-suited for such auditing tasks due to automation bias and opacity (Parasuraman2010), motivating the need for systems that explicitly map observed states to epistemically justifiable actions.
In this work, we use elite-level soccer as a demanding natural laboratory to operationalize this prescriptive auditing paradigm. Substitution decisions share structural properties with many safety-critical contexts: they are time-sensitive, irreversible, and made under pervasive uncertainty. Existing approaches in this domain are limited either by insufficient temporal resolution—introducing exposure bias by favoring cumulative playing time over efficiency (Pappalardo2019; SchmidtLilloBustos2024)—or by a reliance on purely predictive modeling that fails to capture normative tactical necessity (wu; pamukkale).
To enable systematic auditing of human judgment, we propose a hybrid statistical–symbolic framework that overlays intrinsically interpretable fuzzy reasoning onto robust statistical signals. Rather than automating decisions, the system functions as an auditing layer, surfacing decision-relevant risk—such as performance decay or defensive liability—before it manifests as observable failure. Unlike descriptive fuzzy models (Marliere2017; zhou) or static ranking systems (salabun), the proposed approach operationalizes continuous, role-aware evaluation of decision states.
This work makes three contributions. First, we formalize a prescriptive auditing framework for evaluating human decision-making under uncertainty, in which symbolic reasoning contextualizes statistical evidence rather than replacing human agency. Second, we introduce a role-aware cumulative mean metric that eliminates play-time exposure bias, enabling the detection of intra-match performance deterioration. Third, we provide empirical validation in a high-stakes adversarial setting, demonstrating that the proposed auditing system reveals latent risk patterns systematically overlooked by both human experts and black-box predictive models.
2 Results
To evaluate the utility of prescriptive auditing in high-stakes environments, we contrast the proposed Fuzzy Control System (FCS) against state-of-the-art predictive baselines and subject it to a rigorous stress-test using real-world data from elite soccer.
2.1 Overcoming the Predictive Ceiling: Prescription vs. Prediction
Recent benchmarks in soccer analytics reveal a “predictive ceiling” for substitution modeling. mohandas analyzed 51,738 substitutions using supervised learning algorithms (Random Forest, SVM), achieving a maximum accuracy of $\approx 70\%$. This plateau suggests that nearly 30% of tactical decisions are driven by contextual nuances—such as fatigue thresholds or specific game plans—that remain opaque to black-box models.
Critically, supervised models in this domain suffer from a normative flaw: they are trained to mimic the human expert. If historical data encodes risk aversion or delayed reaction times, the ML model learns to replicate these biases. As detailed in Table 1, our proposed framework shifts the objective from minimizing prediction error to maximizing tactical utility, functioning as an auditor rather than an imitator.
Table 1: Structural Comparison: Supervised ML vs. Prescriptive Fuzzy System
Dimension Supervised ML Baseline (e.g., mohandas) Proposed Prescriptive System
Objective Minimize Prediction Error (Mimic Human) Maximize Tactical Utility (Audit Human)
Handling Uncertainty Probabilistic (Confidence Scores) Possibilistic (Membership Degrees)
Interpretability Low (Feature Importance/Black Box) High (Linguistic Rules)
Outcome Bias High (Dependent on realized results) Decoupled (State-based evaluation)
2.2 The Mathematical Mechanism: Eliminating Exposure Bias
A fundamental prerequisite for effective auditing is the ability to detect performance deterioration in real-time. Traditional metrics typically utilize cumulative sum formulations, where a player’s score monotonically increases with playing time regardless of minute-by-minute efficiency. Mathematically, these additive models act as low-pass filters, masking declining performance because a fatigued player performing poorly in the $80^{th}$ minute may still possess a higher total score than a high-impact substitute.
To resolve this Metric Exposure Bias, we reformulated the evaluation metric from a cumulative sum to a Cumulative Mean with Role-Aware Normalization ($P_{cum}$). By calculating the average contribution up to time $t$ and normalizing it against the historical distribution of the agent’s specific tactical role, our metric functions as a high-pass audit. This transformation allows the performance curve to reflect true temporal variability, exposing negative momentum and performance drops that are mathematically invisible in additive models. This methodological shift is what enables the system to detect the latent risks detailed in the following case studies.
2.3 Stress-Testing Decision Latency: The Brazil vs. Belgium Case
We applied the framework to the high-stakes elimination match between Brazil and Belgium (2018 FIFA World Cup) to observe decision latency and normative disagreement. The system outputs ($P_{final}$) were generated in 5-minute intervals.
Figure 1: Temporal evolution of Substitution Priority for Brazil. High values indicate critical inefficacy or risk. Note the divergence in the case of Fagner (Red Dashed Line) compared to executed substitutions.
2.3.1 Alignment and Early Detection
The system successfully captured decision-relevant deterioration prior to human intervention (Table 2).
  • Willian (45’): Priority $P_{final}=72.0$, ranked 4th. Rule R04 (Rapid Drop) flagged a sharp decline in technical actions in the final 20 minutes of the first half. Immediate substitution by the coach confirms offensive ineffectiveness, supported by media reports ESPN_2018.
  • Gabriel Jesus (58’): Priority $P_{final}=99.1$, 2nd highest. Rule R08 (Ineffective Forward) triggered by low cumulative score and no goals/assists. Post-match data: 12 touches in 58 minutes Epoca_2018, validating tactical isolation.
  • Paulinho (73’): Priority $P_{final}=93.1$. System detected ”Early Fatigue” and low participation. External analysis noted minimal offensive impact Gazeta_2018.
2.3.2 The “Fagner Paradox”: Counterfactual Auditing
The most significant finding is the disagreement regarding right-back Fagner. While the human decision-maker retained the player, the system assigned him Maximum Priority (100.0) from minute 45 to 85. This was not a prediction error, but a risk audit. The system aggregated consistently low technical performance with high defensive exposure against opponent Eden Hazard. The realized outcome—Fagner struggling in duels and receiving a yellow card at the 90th minute—validates the system’s “Critical Risk” diagnosis. A predictive ML model trained on this coach’s history would likely have predicted “No Substitution” (True Positive), thereby reinforcing the status quo bias that led to the vulnerability.
Table 2: Model Audit vs. Human Decision (Brazil vs. Belgium)
Player Analysis Slice($t$) $P_{\text{final}}$ Rank Diagnosis Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward Substituted out
Fagner 45’-85’ 100.0 1st Critical Risk Not Substituted
Paulinho 65’-73’ 93.1 3rd Early Fatigue Substituted out
2.3.3 The “Lukaku Paradox”: Latency and Event Masking
Applying the system to the opponent (Belgium, Figure 2) highlights the impact of outcome bias on human decision-making.
Figure 2: Temporal evolution for Belgium. Lukaku reaches maximum priority (100.0) significantly earlier than his substitution.
Romelu Lukaku provided an assist at the 31st minute. While this single salient event seemingly justified his presence to the human manager, the system detected a severe drop in engagement immediately after.
  • Algorithmic Detection (35’): The system identified Lukaku as the #1 substitution candidate immediately post-assist, due to critically low volume (14 ball touches in 87 minutes).
  • Human Latency (+50 min): The coach waited until the 87th minute to substitute him.
This 50-minute latency demonstrates how prescriptive systems can decouple decision quality from “highlight moments,” ensuring consistent performance evaluation (Table 3).
Table 3: Comparative Analysis of Low-Volume Strikers
Player Context FCS Priority Human Response
G. Jesus (BRA) 12 Touches, 0 G/A High (#2) Substituted (58’)
R. Lukaku (BEL) 14 Touches, 1 Assist Max (#1) Delayed Sub (87’)
2.4 Systemic Validation: Longitudinal Patterns
Extending the analysis beyond single instances, we processed a representative subset of 25 World Cup matches—including the complete Brazilian campaign. This broader scope, comprising N = 9,152 decision windows, reveals structural regularities in how the system audits performance.
Technical Anchors vs. Chronic Inefficacy.
The system consistently distinguished between effective stability and passive stagnation. Key players like Coutinho and Marcelo functioned as “Technical Anchors,” maintaining priority scores near zero regardless of match time. Conversely, the system identified “Chronic Inefficacy” in players like Gabriel Jesus across multiple matches, where priority scores escalated without salient errors, supporting the hypothesis that the system detects latent risks invisible to event-based analysis.
Temporal Gap and Anticipation.
A systematic temporal gap was observed between the system’s “Critical” signal ($P_{final}>90$) and human intervention. In cases like Paulinho (vs. Serbia), the system flagged inefficacy several minutes prior to the substitution, quantifying the decision latency inherent in human processing under pressure.
Post-Entry Validation.
By tracking priority scores after a substitution, the system enables counterfactual assessment. Substitutes like Renato Augusto exhibited immediate monotonic decreases in priority (indicating high impact), whereas others (e.g., Fernandinho vs. Switzerland) maintained high risk scores, objectively characterizing the substitution as ineffective.
2.5 Boundary Conditions: The Limits of Performance Auditing
The system’s epistemic boundaries were rigorously tested by the case of Nacer Chadli (Belgium vs. Brazil), who was substituted at the 83rd minute due to an acute injury. This was the only substitution in the analyzed set not anticipated by the FCS ($P_{final}<50.0$).
Far from a systemic failure, this “false negative” provides crucial validation of the framework’s internal logic. Since the injury was a sudden, exogenous event unrelated to prior performance decay or observable fatigue, the system correctly refrained from flagging the player. This confirms that the high priority scores assigned in other cases (e.g., Lukaku, Jesus) were driven by genuine performance degradation, not by a generalized bias toward late-game substitutions. The framework effectively distinguishes between tactical necessity (predictable via data) and force majeure (epistemically inaccessible).
3 Discussion
This study leverages elite sports as a naturalistic laboratory to audit human judgment under uncertainty, revealing systematic divergences between normative risk assessment and expert human behavior. Our primary contribution to human decision science is the quantification of decision latency—the temporal gap between the statistical accumulation of risk and the human reaction to it. By deploying a system designed to decouple decision quality from stochastic outcomes, we provide empirical evidence that human experts in high-stakes environments are subject to structural cognitive biases—specifically status quo bias and outcome bias—which systematically delay necessary interventions.
To rigorize this auditing approach, we formalize the concept of Prescriptive AI:
Definition. Prescriptive AI is a class of intelligent systems whose primary function is to audit, justify, and support human decisions under uncertainty, rather than to automate actions or predict future outcomes.
Unlike standard predictive analytics, this paradigm is defined by five normative objectives:
  • A primary function of decision auditing rather than decision automation;
  • Explicit reasoning under uncertainty and asymmetric risk;
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature (Rudin2019);
  • Contestable recommendations that preserve human agency;
  • Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias (BaronHershey1988).
These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. While the framework itself is model-agnostic, the system presented here serves as a concrete instantiation of these principles. By applying this approach, we show that the limitation of human experts is not a failure of strategic intent, but a failure of real-time information processing: humans struggle to detect non-salient performance decay (the “boiling frog” effect) until a catastrophic or highly salient event occurs.
3.1 Behavioral Audit: Quantifying Cognitive Biases
The empirical disparities between the prescriptive system and human agents are not random errors, but operationalizations of well-documented cognitive heuristics:
The “Fagner Paradox” as Status Quo Bias.
The disagreement regarding the defensive player Fagner illustrates the status quo bias (or omission bias) under pressure. While the system identified a “Critical Risk” state ($P_{final}=100.0$) based on sustained defensive exposure, the human manager chose inaction. This behavior reflects an asymmetric preference for maintaining the current state, requiring substantially stronger evidence to justify change than to justify persistence. Given the informational state available at time $t$, this inertia is normatively difficult to justify.
The “Lukaku Paradox” as Salience Masking and Outcome Bias.
The case of the opposing striker demonstrates how outcome bias distorts real-time evaluation. A single salient positive event (an assist) masked a severe and prolonged deterioration in engagement (over $50$ minutes of inactivity). We term this phenomenon Salience Masking: the overweighting of vivid, recent outcomes at the expense of cumulative base-rate information. The prescriptive system—immune to the “halo effect” of the assist—identified the decision to maintain the player as a high-risk error well before the human agent reacted.
3.2 Implications for Human Decision-Making
These findings extend classical accounts of Bounded Rationality in dynamic environments. We argue that decision latency is not merely a domain-specific artifact, but a domain-general construct reflecting the cognitive cost required to override heuristic inertia in the presence of uncertainty and time pressure.
  • Filtering and Blind Spots: Real-time tactical management requires processing high-dimensional, noisy data streams. To preserve strategic coherence, human agents apply low-pass cognitive filters that attenuate short-term fluctuations. Our results indicate that this adaptive filtering introduces a structural blind spot for gradual but systematic performance decay.
  • AI as Cognitive Orthotics: Accordingly, the role of Prescriptive AI shifts from an “oracle” that predicts future states to a cognitive orthotic—a support structure designed to compensate for specific, predictable human limitations. By surfacing latent risk through interpretable signals (e.g., Rule R04: Rapid Drop), the system confronts decision-makers with evidence their heuristic reasoning would otherwise suppress.
3.3 Normative Foundations: Decision-Making as Epistemic State Transition
The validity of this auditing approach rests on a dynamic conception of rationality, grounded in the logical dynamics framework of vanBenthem2011. Under this view, rationality is a property of the epistemic update at the moment of commitment ($t$), rather than of the stochastic realization of outcomes at a later time ($t+n$). If decision quality is defined prior to outcome realization, then systems designed to support human judgment must operate strictly on the informational state available at the moment of action. This perspective justifies our rejection of outcome-based evaluation metrics: a decision is not “correct” because the team won, but because it constituted a normatively coherent response to the risk state at time $t$.
Naturally, any prescriptive audit is contingent on the chosen normative model of risk. Alternative normative assumptions would yield different—but still auditable—recommendations, without undermining the core principle of decoupling decision quality from outcome realization.
3.4 Ethical Accountability and Epistemic Trust
Finally, we argue that in high-stakes human–AI teaming, interpretability is a pre-condition for legitimacy. Auditing human judgment requires more than predictive accuracy; it requires explanations that enable decision-makers to validate machine assessments against their own reasoning.
Resolving Epistemic Opacity.
In domains subject to intense public scrutiny—such as elite sports, clinical triage, or emergency operations—opaque “black-box” models pose a substantial epistemic risk. When an algorithm recommends a counter-intuitive action (e.g., substituting a star player), a lack of transparency forces stakeholders to choose between blind faith and outright rejection. By providing a transparent audit trail in which each recommendation is traceable to explicit linguistic rules, the proposed fuzzy framework reduces epistemic asymmetry and transforms the interaction from blind obedience into informed deliberation.
The Audit Trail as Institutional Defense.
Beyond decision support, prescriptive systems serve a protective institutional function. In environments vulnerable to suspicion—whether due to financial incentives, reputational risk, or public accountability—a documented audit trail allows decision-makers to demonstrate that their actions (or inaction) were consistent with a normative risk assessment given the information available at time $t$. In this sense, the system functions not merely as a tactical aid, but as an objective witness to the rationality of the decision process.
Ultimately, this work advances prescriptiveness as an ethical design principle. In contexts where decisions are public, consequential, and irreversible, only transparent, auditable, and human-centered systems can achieve the level of trust required to augment—rather than replace—human agency.
4 Methodology
4.1 Baseline Limitations and Methodological Motivation
The academic literature on soccer analytics provides the theoretical foundation for this work, situated at the intersection of performance evaluation and tactical decision modeling.
@@ 4.1 Baseline Limitations and Methodological Motivation
A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo2019, defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
@@
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by mohandas, for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
Finally, the choice of an intrinsically interpretable methodology is not merely a design preference, but a methodological necessity in high-stakes decision-making contexts. As explicitly argued by Rudin2019, post-hoc explanations of black-box models are fundamentally inadequate when decisions have significant consequences, as they obscure the true decision logic and may provide misleading justifications. In domains such as elite soccer, where substitution decisions carry substantial sporting and financial impact, decision support systems must expose their reasoning process in a transparent and auditable manner. This perspective reinforces the limitations of purely predictive machine learning approaches discussed above and provides a principled justification for abandoning black-box optimization in favor of symbolic reasoning. Accordingly, the present work adopts fuzzy logic as the core inference mechanism—rather than as an explanatory add-on—ensuring that substitution recommendations are interpretable, accountable, and directly aligned with domain knowledge and human tactical judgment (Rudin2019).
This perspective finds foundational support in the framework of logical dynamics proposed by vanBenthem2011, which conceptualizes rational agency as a continuous process of information update rather than static optimization. In this context, player substitutions serve as a particularly clear instance of epistemic actions (vanBenthem2007): they do not merely alter the physical parameters of the match, but fundamentally restructure the space of plausible future game states under severe informational constraints. At the moment of decision, the coach operates at the boundary between “hard facts” (e.g., scoreline, remaining time, substitution limits) and “soft information” (e.g., perceived fatigue, tactical momentum). Once executed, a substitution irreversibly fixes certain constraints while reweighting the plausibilities regarding future play. Consequently, decision quality cannot be assessed solely by ex-post outcomes, but by whether the action constituted a rational epistemic update given the information available at time $t$. This framing aligns the proposed system with a dynamic view of rationality—where correctness is understood as the capacity for informed correction under uncertainty—establishing its role not as a predictor of future success, but as an auditor of the epistemic transition induced by the substitution.
Crucially, the proposed approach is hybrid in a functional rather than predictive sense. The statistical components of the system are not employed to forecast future match outcomes or to optimize predictive accuracy. Instead, they serve exclusively to construct a temporally normalized, role-aware representation of the current decision state. The symbolic layer does not refine or correct predictions; it operates as an external normative auditor, evaluating whether continued action is justified given the evidence available at time $t$. In this architecture, hybridization exists to support interpretability, decision auditing, and accountability—not predictive performance.